OpenStudy (anonymous):

Solve 5x + 2y = 7 10x + 4y = 14 a. (1, 1) b. (2,-3/2 ) c. No Solution d. Infinitely Many Solutions

sam (.sam.):

5x + 2y = 7 10x + 4y = 14 -------------------- x = (7-2y)/5 ------------------- 10x + 4y = 14 10(7-2y)/5 + 4y = 14 2(7-2y)+4y=14 14-4y+4y=14 14=14 ----------------- No solution

OpenStudy (lgbasallote):

ahh the good old systems... multiply the first equation by -2... 5(-2)x + 2(-2)y =7(-2) 10x + 4y = 14 -10x -4y = -14 10x + 4y = 14 do you see something odd? :DDD